# Textual Knowledge Matters: Cross-Modality Co-Teaching

### v1

@@ Abstract
In this paper, we study the problem of Generalized Category Discovery (GCD), which aims to cluster unlabeled data from both known and unknown categories using the knowledge of labeled data from known categories. Current GCD methods rely on only visual cues, which however neglect the multi-modality perceptive nature of human cognitive processes in discovering novel visual categories. To address this, we propose a two-phase TextGCD framework to accomplish multi-modality GCD by exploiting powerful Visual-Language Models. TextGCD mainly includes a retrieval-based text generation (RTG) phase and a cross-modality co-teaching (CCT) phase. First, RTG constructs a visual lexicon using category tags from diverse datasets and attributes from Large Language Models, generating descriptive texts for images in a retrieval manner. Second, CCT leverages disparities between textual and visual modalities to foster mutual learning, thereby enhancing visual GCD. In addition, we design an adaptive class aligning strategy to ensure the alignment of category perceptions between modalities as well as a soft-voting mechanism to integrate multi-modality cues. Experiments on eight datasets show the large superiority of our approach over state-of-the-art methods. Notably, our approach outperforms the best competitor, by 7.7% and 10.8% in All accuracy on ImageNet-1k and CUB, respectively.
@@ 1 Introduction
Despite the remarkable advancements of deep learning in visual recognition, a notable criticism is that the models, once trained, show a significant limitation in recognizing novel classes not encountered during the supervised training phase. Drawing inspiration from the innate human capacity to seamlessly acquire new knowledge with reference to previously assimilated information, Generalized Category Discovery (GCD) Vaze et al. (2022) is proposed to leverage the knowledge of labeled data from known categories to automatically cluster unlabeled data that belong to both known and unknown categories. While current GCD methods Vaze et al. (2022); Wen et al. (2023); Pu et al. (2023); Zhao et al. (2023); Zhang et al. (2023) have demonstrated considerable success utilizing advanced large-scale visual models (e.g., ViT Dosovitskiy et al. (2020)), as illustrated in Fig. 1, they predominantly focus on visual cues. In contrast, human cognitive processes for identifying novel visual categories usually incorporate multiple modalities Sloutsky (2010), such as encompassing visual, auditory, and textual elements in recognizing a subject. In the light of this, unlike existing GCD methods Vaze et al. (2022); Wen et al. (2023); Pu et al. (2023); Zhao et al. (2023); Zhang et al. (2023); Wang et al. (2023) that rely on only visual modality, we propose to exploit both visual and textual cues for GCD.
In this paper, we advocate the utilization of large-scale pre-trained Visual-Language Models (VLMs) Radford et al. (2021); Yu et al. (2022); Li et al. (2023b) to inject rich textual information into GCD. However, VLMs require predetermined, informative textual descriptors (e.g., class names) for matching or recognizing images. This poses a significant challenge in GCD involving unlabeled data, particularly for unknown categories that lack predefined class names. Hence, how to provide relevant textual cues for unlabeled data is of a key in facilitating GCD with textual information.
@@
To this end, we introduce the TextGCD framework, comprising a retrieval-based text generation (RTG) phase and a cross-modality co-teaching (CCT) phase, to fully leverage the mutual benefit of visual and textual cues for solving GCD. In particular, RTG initially constructs a visual lexicon encompassing category tags from diverse datasets and attributes gleaned from Large Language Models (LLMs), which offer categorical insights. Subsequently, the most pertinent tags and attributes are selected from the lexicon to craft descriptive texts for all images through an offline procedure. While the augmentation of textual information for unlabeled images via RTG is beneficial, the effective integration of this textual information to enhance visual category discovery remains an essential challenge.
To solve the challenge, we introduce CCT to leverage the inherent disparities between textual and visual modalities, harnessing these distinctions to facilitate joint learning, which is important for co-teaching Han et al. (2018). To this end, we devise a cross-modal co-teaching training strategy that fosters mutual learning and collective enhancement between text and image models. Additionally, to ensure the effective alignment of category perceptions between text and image models and to facilitate comprehensive learning from both modalities, we introduce a warm-up stage and a class-aligning stage before engaging in cross-modality co-teaching. Finally, a soft-voting mechanism is deployed to amalgamate insights from both modalities, aiming to enhance the accuracy of category determination. In short, introducing the textual modality and exchanging insights between modalities have significantly improved the precision of visual category identification, particularly for previously unseen categories. Our contributions are summarized as follows:
@@ 2 Related Work
Generalized Category Discovery (GCD) aims to accurately classify an unlabeled set containing both known and unknown categories, based on another dataset labeled with only known categories. Vaze et al. Vaze et al. (2022) first proposed optimizing image feature similarities through supervised and self-supervised contrastive learning to address GCD. Building on this, SimGCD Wen et al. (2023) designed a parametric classification baseline for GCD. For better image representations, DCCL Pu et al. (2023) proposed a novel approach to contrastive learning at both the concept and instance levels. Similarly focusing on image representation, PromptCAL Zhang et al. (2023) introduced a Contrastive Affinity Learning approach with auxiliary visual prompts designed to amplify the semantic discriminative power of the pre-trained backbone. Furthermore, SPTNet Wang et al. (2023) proposed iteratively implementing model finetuning and prompt learning, resulting in clearer boundaries between different semantic categories. Unlike existing methods that primarily focus on visual cues, we design a retrieval-based approach to introduce text cues from LLM for unknown categories. Recently, CLIP-GCD Ouldnoughi et al. (2023) concatenated image features with textual features obtained from a Knowledge Database and categorized them using a clustering method. Unlike CLIP-GCD, which merely concatenates text and visual features from a frozen backbone, we leverage the differences between textual and visual models to establish a dynamic co-teaching scheme.
@@
Visual-Language Models (VLMs) are designed to map images and text into a unified embedding space, facilitating cross-modal alignment. Among the prominent works, CLIP Radford et al. (2021) employs contrastive representation learning with extensive image-text pairs, showcasing remarkable zero-shot transfer capabilities across a variety of downstream tasks. Additionally, LENS Berrios et al. (2023) leverages VLMs as visual reasoning modules, integrating them with Large Language Models (LLMs) for diverse visual applications. As VLMs effectively bridge the visual and textual modalities, enhancing image classification tasks with knowledge from LLMs has been extensively explored Menon & Vondrick (2022); Yang et al. (2023); Novack et al. (2023). However, these works use VLMs to enhance image classification in scenarios where the names of all classes are predefined for VLMs. In this paper, we exploit VLMs to address the GCD task, where the unlabeled data do not have exploitable textual information for VLMs. We utilize a retrieval-based text generation method to enable the GCD task to benefit from large VLMs.
@@
Co-teaching was originally proposed as a learning paradigm to address the issue of label noise Han et al. (2018). This strategy involves training two peer networks that select low-loss instances from a noisy mini-batch to train each other. Nested-Co-teaching Chen et al. (2021) and JoCoR Wei et al. (2020) further introduce compression regularization and co-regularization for each training example to optimize the co-teaching process. Yuan et al. Yuan et al. (2024) established three symmetric peer proxies with pseudo-label-driven co-teaching to address the Offline model-based optimization task. Yang et al. Yang et al. (2020) were the first to introduce the co-teaching strategy into the Person re-identification task, designing an asymmetric co-teaching framework. Similarly, Roy et al. Roy et al. (2021) pioneered co-teaching between classifiers of a dual classifier head to tackle the multi-target domain task. Unlike these works, our focus is on applying co-teaching to the GCD task. Furthermore, rather than employing two peer networks within the same modality or learning from self-discrepancy, we initiate co-teaching between the image model and the text model, leveraging the disparities between modalities to foster mutual enhancement.
@@ 3 Method
Task Configuration: In GCD, we are given a labeled dataset, denoted as $\mathcal{D}_{L}=\{(\mathbf{x}_{i},y_{i}^{l})\}_{i=1}^{M}\subseteq\mathcal{X}\times\mathcal{Y}_{L}$, and an unlabeled dataset, denoted as $\mathcal{D}_{U}=\{(\mathbf{x}_{i},y_{i}^{u})\}_{i=1}^{N}\subseteq\mathcal{X}\times\mathcal{Y}_{U}$, where $M$ and $N$ indicate the number of samples in the $\mathcal{D}_{L}$ and $\mathcal{D}_{U}$, respectively. $\mathcal{Y}_{L}$ and $\mathcal{Y}_{U}$ indicate the label spaces for labeled and unlabeled datasets, respectively, where $\mathcal{Y}_{L}\subseteq\mathcal{Y}_{U}$. That is, $\mathcal{D}_{U}$ contains data from unknown categories, and $\mathcal{Y}_{U}$ is unavailable. The objective of GCD is to leverage the prior knowledge from known categories within $\mathcal{D}_{L}$ to classify samples in $\mathcal{D}_{U}$ effectively. Following Wen et al. (2023), it is assumed that the number of classes in $\mathcal{D}_{U}$, represented by $K=|{\mathcal{Y}}_{U}|$, is predetermined.
@@ 3.2 Baseline
We follow the SimGCD Wen et al. (2023) to build our parametric learning baseline, which consists of a supervised loss for labeled data and an unsupervised loss for both labeled and unlabeled data.
Specifically, the supervised loss utilizes the conventional cross-entropy loss, defined as:
$\mathcal{L}_{\text{sup}}=\frac{1}{|B^{l}|}\sum_{i\in B^{l}}\ell(\bm{y}_{i},\bm{p}_{i}),$ (3.1)
where $B^{l}$ indicates the mini-batch for labeled data, $\ell$ is the traditional cross-entropy loss, and $\bm{p}_{i}=\sigma({\mathcal{H}(\mathcal{E}(\mathbf{x}_{i}))/\tau})$ is the predicted probabilities of input $\mathbf{x}_{i}$. Here, $\mathcal{E}$ represents the backbone encoder, and $\mathcal{H}$ indicates the parametric classifier. $\sigma(\cdot)$ denotes the softmax function, and $\tau$ is a temperature parameter set to $\tau_{s}$.
@@
$\mathcal{L}_{\text{unsup}}=\frac{1}{|B|}\sum_{i\in B}\ell({\bm{q}^{\prime}_{i}},\bm{p}_{i})-\varepsilon H(\overline{\bm{p}}),$ (3.2)
where $\bm{q}^{\prime}_{i}$ is the predicted probabilities of $\mathbf{x}_{i}^{\prime}$ using a sharper temperature value $\tau_{t}$. We also include a mean-entropy maximization regularizer $H(\overline{\bm{p}})$ Assran et al. (2022) for the unsupervised objective. Here, $H(\overline{\bm{p}})$= $-\sum_{k}\overline{\bm{p}}^{(k)}\log\overline{\bm{p}}^{(k)}$, where $\overline{\bm{p}}=\frac{1}{2|B|}\sum_{i\in B}(\bm{p}_{i}+\bm{p}_{i}^{\prime})$. $\bm{p}_{i}$ and $\bm{p}_{i}^{\prime}$ are the probabilities of $\mathbf{x}_{i}$ and $\mathbf{x}_{i}^{\prime}$, respectively, which use the same temperature of $\tau_{u}$. $B$ indicates the mini-batch for both labeled and unlabeled data. The hyperparameter $\varepsilon$ aligns with the configuration used in SimGCD Wen et al. (2023).
@@
$\mathcal{L}_{\text{base}}=\lambda\cdot\mathcal{L}_{\text{sup}}+(1-\lambda)\cdot\mathcal{L}_{\text{unsup}},$ (3.3)
@@ 3.3 Retrieval-based Text Generation
Building the Visual Lexicon. To construct a visual lexicon, we aggregate a diverse and extensive collection of tags from established benchmarks in semantic segmentation, object detection and classification. These benchmarks are selected for their comprehensive coverage of the visual categories in the world. Additionally, we utilize a large language model (e.g., GPT3 Brown et al. (2020)) to further describe each tag by using the prompt of “What are useful features for distinguishing a {tag} in a photo?”, which can augment the visual lexicon with the attribute of “{tag} which has the {feature}”.
Tag Retrieval. Given the visual lexicon, comprising tags and corresponding attributes, we utilize an auxiliary VLM (e.g., CLIP Radford et al. (2021)) to encode it, thereby creating a textual feature bank. For each image, we use the auxiliary model to generate its visual feature, which is then compared against the entries in the textual feature bank using cosine similarity. Subsequently, we identify the most relevant $n_{t}$ tags and $n_{a}$ attributes exhibiting the highest similarity to the image. We then concatenate the textual descriptors of these tags and attributes based on the similarity ranking. This process yields the categorical descriptive text $\mathbf{t}_{i}$ for each image $\mathbf{x}_{i}$. The workflow of the RTG phase is delineated in Fig. 3.
Discussion. In this paper, we aim to harness the capabilities of the existing foundational models to deliver comprehensive and accurate textual information for GCD. To this end, we employ GPT-3 Brown et al. (2020) as the LLM to obtain attributes and the ViT-H-based CLIP Radford et al. (2021) as the auxiliary VLM to identify tags and attributes. We also attempt to use a smaller auxiliary VLM, e.g., ViT-B-based CLIP, to generate the descriptive text. However, it demonstrated limited efficacy for fine-grained classification because ViT-B-based CLIP tends to focus on more general Tags and Attributes in the Visual Lexicon. Using Flower102 for instance, ViT-B-based CLIP tends to give high similarity to broader Tags like “Plant & Flower”. Compared to existing methods that only use visual cues for GCD, we explore the possibility of introducing the textual information to facilitate the GCD task by using the freely available, public foundational models, which would be a new trend in the community. In addition, our approach enjoys several advantages. Firstly, the RTG is an offline phase, where we only need to process each sample once, avoiding the need for excessive computational overhead. Secondly, given the generated descriptive text, we only need to train the GCD model with a smaller model (ViT-B-based CLIP), achieving significantly higher results than existing methods. Third, the auxiliary VLM can be replaced with other advanced foundational models, such as FLIP Li et al. (2023b) and CoCa Yu et al. (2022). Consequently, the effectiveness of our method is likely to be enhanced in line with the advancements in these foundational models.

### v2

@@ Abstract
In this paper, we study the problem of Generalized Category Discovery (GCD), which aims to cluster unlabeled data from both known and unknown categories using the knowledge of labeled data from known categories. Current GCD methods rely on only visual cues, which however neglect the multi-modality perceptive nature of human cognitive processes in discovering novel visual categories. To address this, we propose a two-phase TextGCD framework to accomplish multi-modality GCD by exploiting powerful Visual-Language Models. TextGCD mainly includes a retrieval-based text generation (RTG) phase and a cross-modality co-teaching (CCT) phase. First, RTG constructs a visual lexicon using category tags from diverse datasets and attributes from Large Language Models, generating descriptive texts for images in a retrieval manner. Second, CCT leverages disparities between textual and visual modalities to foster mutual learning, thereby enhancing visual GCD. In addition, we design an adaptive class aligning strategy to ensure the alignment of category perceptions between modalities as well as a soft-voting mechanism to integrate multi-modality cues. Experiments on eight datasets show the large superiority of our approach over state-of-the-art methods. Notably, our approach outperforms the best competitor, by 7.7% and 10.8% in All accuracy on ImageNet-1k and CUB, respectively. Code is available at https://github.com/HaiyangZheng/TextGCD.
Keywords:
Generalized Category Discovery Cross-Modality Co-Teaching
@@ 1 Introduction
Despite the remarkable advancements of deep learning in visual recognition, a notable criticism is that the models, once trained, show a significant limitation in recognizing novel classes not encountered during the supervised training phase. Drawing inspiration from the innate human capacity to seamlessly acquire new knowledge with reference to previously assimilated information, Generalized Category Discovery (GCD) [40] is proposed to leverage the knowledge of labeled data from known categories to automatically cluster unlabeled data that belong to both known and unknown categories. While current GCD methods [40, 43, 32, 51, 49] have demonstrated considerable success utilizing advanced large-scale visual models (e.g., ViT [9]), as shown in Fig. 1, they predominantly focus on visual cues. In contrast, human cognitive processes for identifying novel visual categories usually incorporate multiple modalities [38], such as encompassing visual, auditory, and textual elements in recognizing a subject. In the light of this, unlike existing GCD methods [40, 43, 32, 51, 49, 42] that rely on only visual modality, we propose to exploit both visual and textual cues for GCD.
In this paper, we advocate the utilization of large-scale pre-trained Visual-Language Models (VLMs) (i.e. CLIP [33]) to inject rich textual information into GCD. However, VLMs require predetermined, informative textual descriptors (e.g., class names) for matching or recognizing images. This poses a significant challenge in GCD involving unlabeled data, particularly for unknown categories that lack predefined class names. Hence, how to provide relevant textual cues for unlabeled data is of a key in facilitating GCD with textual information.
@@
To this end, we introduce the TextGCD framework, comprising a retrieval-based text generation (RTG) phase and a cross-modality co-teaching (CCT) phase, to leverage the mutual benefit of visual and textual cues for solving GCD. In particular, RTG constructs a visual lexicon encompassing category tags from diverse datasets and attributes gleaned from Large Language Models (LLMs), which offer categorical insights. Subsequently, the most pertinent tags and attributes are selected from the lexicon to craft descriptive texts for all images through an offline procedure. While the augmentation of textual information for unlabeled images via RTG is beneficial, the effective integration of this textual information to enhance visual category discovery remains an essential challenge.
To solve the challenge, we introduce CCT to leverage the inherent disparities between textual and visual modalities, harnessing these distinctions to facilitate joint learning, which is important for co-teaching [12]. To this end, we devise a cross-modal co-teaching training strategy that fosters mutual learning and collective enhancement between text and image models. Additionally, to ensure the effective alignment of category perceptions between text and image models and to facilitate comprehensive learning from both modalities, we introduce a warm-up stage and a class-aligning stage before engaging in cross-modality co-teaching. Finally, a soft-voting mechanism is deployed to amalgamate insights from both modalities, aiming to enhance the accuracy of category determination. In short, introducing the textual modality and exchanging insights between modalities have significantly improved the precision of visual category identification, particularly for previously unseen categories. We make the following contributions:
@@ 2 Related Works
Generalized Category Discovery (GCD) aims to accurately classify an unlabeled set containing both known and unknown categories, based on another dataset labeled with only known categories. It is an extension of novel category discovery [53, 36, 13, 14, 10, 52] where the unlabeled set only contains unknown classes. Vaze et al. [40] first proposed optimizing image feature similarities through supervised and self-supervised contrastive learning to address GCD. Building on this, SimGCD [43] designed a parametric classification baseline for GCD. For better image representations, DCCL [32] proposed a novel approach to contrastive learning at both the concept and instance levels. Similarly focusing on image representation, PromptCAL [49] introduced a Contrastive Affinity Learning approach with auxiliary visual prompts designed to amplify the semantic discriminative power of the pre-trained backbone. Furthermore, SPTNet [42] proposed iteratively implementing model finetuning and prompt learning, resulting in clearer boundaries between different semantic categories. Unlike existing methods that primarily focus on visual cues, we design a retrieval-based approach to introduce text cues from LLM for unknown categories. Recently, CLIP-GCD [30] concatenated image features with textual features obtained from a Knowledge Database and categorized them using a clustering method. Unlike CLIP-GCD, which merely concatenates text and visual features from a frozen backbone, we leverage the differences between textual and visual models to establish a dynamic co-teaching scheme.
@@
Visual-Language Models (VLMs) are designed to map images and text into a unified embedding space, facilitating cross-modal alignment. Among the prominent works, CLIP [33] employs contrastive representation learning with extensive image-text pairs, showcasing remarkable zero-shot transfer capabilities across a variety of downstream tasks. Additionally, LENS [2] leverages VLMs as visual reasoning modules, integrating them with Large Language Models (LLMs) for diverse visual applications. As VLMs effectively bridge the visual and textual modalities, enhancing image classification tasks with knowledge from LLMs has been extensively explored [27, 46, 29]. However, these works use VLMs to enhance image classification in scenarios where the names of all classes are predefined for VLMs. In this paper, we exploit VLMs to address the GCD, where the unlabeled data do not have exploitable textual information for VLMs. We utilize a retrieval-based text generation method to enable the GCD to benefit from VLMs.
@@
Co-teaching, which stemmed from the Co-training approach [3], was originally proposed as a learning paradigm to address the issue of label noise, [12]. This strategy involves training two peer networks that select low-loss instances from a noisy mini-batch to train each other. Yuan et al. [48] established three symmetric peer proxies with pseudo-label-driven co-teaching to address the Offline model-based optimization task. Yang et al. [45] were the first to introduce the co-teaching strategy into the Person re-identification task, designing an asymmetric co-teaching framework. Similarly, Roy et al. [35] pioneered co-teaching between classifiers of a dual classifier head to tackle the multi-target domain task. Unlike these works, our focus is on applying co-teaching to the GCD task. Furthermore, rather than employing two peer networks within the same modality or learning from self-discrepancy, we initiate co-teaching between the image model and the text model, leveraging the disparities between modalities to foster mutual enhancement.
@@ 3 Method
Task Configuration: In GCD, we are given a labeled dataset, denoted as $\mathcal{D}_{L}=\{(\mathbf{x}_{i},y_{i}^{l})\}_{i=1}^{M}\subseteq\mathcal{X}\times\mathcal{Y}_{L}$, and an unlabeled dataset, denoted as $\mathcal{D}_{U}=\{(\mathbf{x}_{i},y_{i}^{u})\}_{i=1}^{N}\subseteq\mathcal{X}\times\mathcal{Y}_{U}$, where $M$ and $N$ indicate the number of samples in the $\mathcal{D}_{L}$ and $\mathcal{D}_{U}$, respectively. $\mathcal{Y}_{L}$ and $\mathcal{Y}_{U}$ indicate the label spaces for labeled and unlabeled datasets, respectively, where $\mathcal{Y}_{L}\subseteq\mathcal{Y}_{U}$. That is, $\mathcal{D}_{U}$ contains data from unknown categories, and $\mathcal{Y}_{U}$ is unavailable. The objective of GCD is to leverage the prior knowledge from known categories within $\mathcal{D}_{L}$ to classify samples in $\mathcal{D}_{U}$ effectively. Following [43], it is assumed that the number of classes in $\mathcal{D}_{U}$, represented by $K=|{\mathcal{Y}}_{U}|$, is predetermined.
@@ 3.2 Baseline
We follow the SimGCD [43] to build our parametric learning baseline, which consists of a supervised loss for labeled data and an unsupervised loss for both labeled and unlabeled data.
Specifically, the supervised loss utilizes the conventional cross-entropy loss:
$\mathcal{L}_{\text{sup}}=\frac{1}{|B^{l}|}\sum_{i\in B^{l}}\ell(\boldsymbol{y}_{i},\boldsymbol{p}_{i}),$ (1)
where $B^{l}$ indicates the mini-batch for labeled data, $\ell$ is the cross-entropy loss, and $\boldsymbol{p}_{i}=\sigma({\mathcal{H}(\mathcal{E}(\mathbf{x}_{i}))/\tau})$ is the predicted probabilities of input $\mathbf{x}_{i}$. Here, $\mathcal{E}$ represents the backbone encoder, and $\mathcal{H}$ indicates the parametric classifier. $\sigma(\cdot)$ denotes the softmax function, and $\tau$ is a temperature parameter set to $\tau_{s}$.
@@
$\mathcal{L}_{\text{unsup}}=\frac{1}{|B|}\sum_{i\in B}\ell({\boldsymbol{q}^{\prime}_{i}},\boldsymbol{p}_{i})-\varepsilon H(\overline{\boldsymbol{p}}),$ (2)
where $\boldsymbol{q}^{\prime}_{i}$ is the predicted probabilities of $\mathbf{x}_{i}^{\prime}$ using a sharper temperature value $\tau_{t}$. We also include a mean-entropy maximization regularizer $H(\overline{\boldsymbol{p}})$ [1] for the unsupervised objective. $H(\overline{\boldsymbol{p}})$=$-\sum_{k}\overline{\boldsymbol{p}}^{(k)}\log\overline{\boldsymbol{p}}^{(k)}$, where $\overline{\boldsymbol{p}}=\frac{1}{2|B|}\sum_{i\in B}(\boldsymbol{p}_{i}+\boldsymbol{p}_{i}^{\prime})$. $\boldsymbol{p}_{i}$ and $\boldsymbol{p}_{i}^{\prime}$ are the probabilities of $\mathbf{x}_{i}$ and $\mathbf{x}_{i}^{\prime}$, respectively, which use the same temperature of $\tau_{u}$. $B$ indicates the mini-batch for both labeled and unlabeled data. The hyperparameter $\varepsilon$ aligns with the configuration used in SimGCD [43].
@@
$\mathcal{L}_{\text{base}}=\lambda\cdot\mathcal{L}_{\text{sup}}+(1-\lambda)\cdot\mathcal{L}_{\text{unsup}},$ (3)
@@ 3.3 Retrieval-based Text Generation
Building the Visual Lexicon. To construct a visual lexicon, we aggregate a diverse and extensive collection of tags from established benchmarks in semantic segmentation, object detection and classification. These benchmarks are selected for their comprehensive coverage of the visual categories in the world. Additionally, we utilize a large language model (e.g., GPT3 [5]) to further describe each tag by using the prompt of “What are useful features for distinguishing a {tag} in a photo?”, which can augment the visual lexicon with the attribute of “{tag} which has the {feature}”.
Tag Retrieval. Given the visual lexicon, comprising tags and corresponding attributes, we utilize an auxiliary VLM (e.g., CLIP [33]) to encode it, thereby creating a textual feature bank. For each image, we use the auxiliary model to generate its visual feature, which is then compared against the entries in the textual feature bank using cosine similarity. Subsequently, we identify the most relevant $n_{t}$ tags and $n_{a}$ attributes exhibiting the highest similarity to the image. We then concatenate the textual descriptors of these tags and attributes based on the similarity ranking. This process yields the categorical descriptive text $\mathbf{t}_{i}$ for each image $\mathbf{x}_{i}$. The workflow of the RTG phase is delineated in Fig. 3.
Discussion. In this paper, we aim to harness the capabilities of the existing foundational models to deliver comprehensive and accurate textual information for GCD. To this end, we employ GPT-3 [5] as the LLM to obtain attributes and the ViT-H-based CLIP [33] as the auxiliary VLM to identify tags and attributes. We also attempt to use a smaller auxiliary VLM, e.g., ViT-B-based CLIP, to generate the descriptive text. However, it demonstrated limited efficacy for fine-grained classification because ViT-B-based CLIP tends to focus on more general Tags and Attributes in the Visual Lexicon. Using Flower102 for instance, ViT-B-based CLIP tends to give high similarity to broader Tags like “Plant & Flower”. Compared to existing methods that only use visual cues for GCD, we explore the possibility of introducing the textual information to facilitate the GCD task by using the freely available, public foundational models, which would be a new trend in the community. In addition, our approach enjoys several advantages. Firstly, the RTG is an offline phase, where we only need to process each sample once, avoiding the need for excessive computational overhead. Secondly, given the generated descriptive text, we only need to train the GCD model with a smaller model (ViT-B-based CLIP), achieving significantly higher results than existing methods. Third, the auxiliary VLM can be replaced with other advanced foundational models, such as FLIP [23] and CoCa [47]. Consequently, the effectiveness of our method is likely to be enhanced in line with the advancements in these foundational models.